# Floating orbits and energy extraction from magnetized Kerr black holes

### v1

@@ Abstract
We study electromagnetic radiation reaction on a charged particle around a weakly magnetized Kerr black hole. We solve numerically the Teukolsky equation to find energy fluxes in electromagnetic radiation at the horizon and at spatial infinity. We also employ analytical methods in the low-frequency limit, finding excellent agreement with the numerical results. For a wide range of parameters, energy fluxes on the horizon are negative for all orbital radii: the modes are amplified through superradiance. More interestingly, the flux on the horizon is larger (in absolute value) than the flux at infinity for orbits with orbital radius $r_{0}\gtrsim 9M$floating orbits are a generic outcome of having magnetic fields around black holes. Particles on these orbits extract energy from the BH through their radiation field; the net effect is an increase in the particles’ kinetic energy, so that they outspiral to infinity. Although in realistic accretion disks viscous forces are expected to dominate, floating orbits remain an interesting feature of the existence of ergoregions in black hole spacetimes.
@@ I Introduction
The Kerr solution to the vacuum Einstein field equations [1] describes the unique stationary and asymptotically flat black hole (BH) solution in General Relativity [2, 3, 4, 5]. It represents an isolated rotating BH, and is characterized by two parameters only – its mass and angular momentum. This is the simplest theoretical description of any macroscopic object known to date!
In the last decade, there has been mounting experimental evidence pointing to BHs being commonplace in the Universe. Namely, we refer to the observations of gravitational waves from binary BH coalescences in the LIGO-Virgo-Kagra observatories [6, 7, 8] and of supermassive BHs at the center of galaxies [9, 10, 11]. So far, all observations are compatible with the dark compact objects being Kerr BHs described by General Relativity. The “Kerr paradigm” [12] will be tested with increasing precision in the near future with space-based [13] and other next generation detectors [14, 15, 16]. Thus, the study of the rich physics occurring in the vicinity of BHs is an active and timely topic.
Astrophysical BHs are not in isolation: instead, they are immersed in complex and dynamic systems [9, 10, 17, 18, 19, 11]. For example, BHs may be surrounded by an accretion disk of ionized matter [17, 18, 19, 20, 21, 22, 23], which can support appreciably strong magnetic fields. Indeed, recent observation show that BHs are commonly immersed in magnetic fields of order $\sim 10^{4}$ and $\sim 10^{8}$ Gauss, for supermassive and stellar mass BHs, respectively [24, 25, 26, 27, 19]. Here, we consider a weak magnetic field, asymptotically uniform and aligned with the BH axis of rotation. We take the magnetic field weak enough that the geometry is described by the Kerr metric [28, 29]. This zero backreaction approximation is valid only if we restrict to a region of spacetime satisfying [30]
@@
Kerr BHs are prone to energy extraction mechanisms [31, 32, 33, 34, 35, 36, 37, 38, 39], which convert the BH rotational energy into kinetic energy of particles or occupation number of fields. In this paper we report an interesting effect of superradiance [35], whereby a charged particle orbiting a weakly magnetized Kerr BH can extract rotational energy from the latter and outspiral. This finding is another example of how magnetic fields can provide means of energy extraction from BHs [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. In a previous paper [49] we studied the the effect of radiation reaction in the weakly magnetized Schwarzschild system. We found a “horizon dominance effect” for a specific orbital configuration – the plus configuration. In this work, we will look at a similar orbital configuration – the prograde plus configuration – and show that the radiation emitted by the particle is now amplified through superradiance. This amplification is strong enough for the particle to experience a net energy gain and outspiral to infinity, following a so-called floating orbit. This type of orbit cannot be found in the unmagnetized Kerr system [50], unless massive fundamental fields are invoked [51].
We focus on charged particles in equatorial-plane circular orbits around weakly magnetized Kerr BHs. We use the Teukolsky formalism [52, 53] to obtain the electromagnetic (EM) energy and angular momentum fluxes on the horizon and at infinity [54, 55, 56, 57]. By energy and angular momentum conservation, knowledge of the fluxes allows one to evolve the orbital parameters and study particle trajectories, as is usually done in the absence of magnetic fields [56, 57, 50, 51, 58, 59]. Although this orbital evolution is not done here, we show in Appendix A that particles in circular orbit do remain in circular orbit when driven by radiation reaction. Note that this calculation will only capture dissipative effects, even though the full self-force also includes a conservative counterpart [60, 61, 62].
@@
We use geometric units $G=c=1$ and a Gaussian system with $14\pi\varepsilon_{0}=\mu_{0}/4\pi=$ 1. The metric signature is (,+,+,+)(-,\,+,\,+,\,+)( - , + , + , + ) and Greek indices run from 00 to $3$.
@@ II.1 Weakly magnetized Kerr BH
Figure 1: Schematic representation of the four configurations of equatorial circular orbits in the weakly magnetized Kerr BH system. The BH is immersed in an asymptotically uniform magnetic field; both the BH angular momentum and the asymptotic magnetic field are orthogonal to the plane of this page. The charged particle (grey dot) is assumed to orbit the BH (black disk at the center) anticlockwise; the BH rotation, parametrized by $a$. is indicated by black arrows in the center; the cyclotron frequency is $\omega_{c}=qB_{0}/m$, where $B_{0}$ is the magnetic field strength at infinity, $q$ is the charge of the particle, and $m$ is its mass. The configurations are invariant under reflections about the plane of the paper. We also include a Newtonian interpretation of the difference between the plus and minus configurations: $\bm{F}_{\text{N}}$ and $\bm{F}_{\text{L}}$ are respectively the gravitational and Lorentz forces; the latter is centripetal (centrifugal) in minus (plus) configuration orbits.
@@
and $0\leq|a|\leq M$. The inner and outer horizons are located at the roots of $\Delta$,
@@ II.2 Circular orbit motion
$H\left(\pi_{\mu},x^{\mu}\right)=\frac{1}{2}\left[g_{rr}\pi_{r}^{2}+g_{rr}\pi_{\theta}^{2}+m^{2}\mathcal{F}\right]\,,$ (8)
@@
This system exhibits a reflection symmetry $(a,\mathcal{L},\omega_{c})\to(-a,-\mathcal{L},-\omega_{c})$, so that circular orbits can be divided into four different configurations. Under the assumption $\mathcal{L}>0$, these are characterized as follows:
@@
We obtained a full characterization of the two configurations present in the case of a magnetized Schwarzschild BH [49], where there is no distinction between prograde and retrograde orbits. The Kerr case is qualitatively similar, the only difference being the orbital frequency of particles following geodesics, which is now given by Eq. (6).
In the study of the magnetized Schwarzschild BH [49], we found a “horizon dominance effect” in plus configuration orbits: for sufficiently large $M\omega_{c}$, the energy flux on the horizon dominates the energy flux at infinity, even for arbitrarily wide circular orbits. This effect occurs for low orbital frequency $M\Omega_{0}\ll 1$, so that the radiation emitted is also of low frequency. Thus, when the BH is rotating, these modes are candidates for superradiant amplification [35]. The purpose of this work is to investigate this possibility, so we focus on the most promising configuration, the prograde plus configuration.
@@
$\Omega_{0}\sim\omega_{+}\quad(r_{0}\ll r_{c})\ \ ,\qquad\Omega_{0}\sim\frac{\omega_{+}^{2}}{\omega_{c}}\quad(r_{0}\ll r_{c})\,.$ (15)
We will only study one particular set of parameters in this work. These are
@@
These are motivated by gravitational wave observations of nearly equal mass mergers [67, 68]. The value of $\omega_{c}$ can be as high as $M\omega_{c}\sim 10^{11}$ for an electron orbiting a supermassive BH, according to recent measurements of the magnetic fields surrounding the latter [24, 25, 26, 27]. We chose a more conservative value due to limited numerical precision. Still, it is important to note that for different values of $a$ and $\omega_{c}$ we found qualitatively similar results.
Onwards, we focus on prograde plus configuration orbits with the system parameters shown of Eq. (16). The orbital velocity profile is shown in Fig. 2 for radii outside the innermost circular orbit (ISCO), which we find to be
@@ III.1 The Teukolsky equation
and similarly for the source $T$. The ${}_{s}S_{\omega\ell}(\theta)$ are the spin weighted spheroidal harmonics [72, 73]. Separation of variables yields two ordinary differential equations, one in the radial coordinate $r$ and another in the $\theta$ coordinate. The angular equation is a regular Sturm-Liouville problem with eigenvalues ${}_{s}E_{\omega\ell}$ and defines spin-weighted spheroidal harmonics. The radial equation is a Schrödinger-like equation; omitting the subscripts in $R$ and $T$ for simplicity it takes the form
@@
Thus, the problem amounts to solving the radial equation (19) with appropriate boundary conditions, corresponding to purely ingoing waves ot the horizon and purely outgoing waves at infinity. This is done by first constructing two linearly independent solutions – $R^{H}$ and $R^{\infty}$ – to the homogeneous problem ($T=0$), such that they satisfy the intended boundary conditions at the horizon and at infinity, respectively. A straightforward calculation [53] reveals that these solutions must have the following asymptotics
@@
The solution to the non-homogeneous problem ($T\neq 0$) is simply
@@ III.2 EM radiation by charged particle in a circular orbit
where $J_{2}$ is defined in Eq. 3.8 of Ref. [53]. The functional form of the EM 4-current $J_{\mu}$ is the same as in the case where the background is the Scharzschild solution and reads [49]
@@
where $\omega=m\Omega_{0}$, $\Theta(\cdot)$ is the Heaviside step function, $g_{\mu\nu}$ are the components of the Kerr metric (2), and all quantities in $Z^{\infty,H}_{\ell m}$ are evaluated at $r=r_{0}$, $\theta=\pi/2$; recall that the functions $R^{\infty,H}$ carry $\omega\ell m$ subscripts.
@@
Thus, once the homogeneous radial Eq. (19) is solved, we can determine the full solution. With $R^{\infty,H}$ we can determine the energy and angular momentum fluxes in a straightforward manner. We will solve the homogeneous equation by employing both analytical and numerical methods.
@@ III.3 Comment on adiabatic evolution
Our results can be used to study the motion of a charged particle around a weakly magnetized Kerr BH, by taking a sequence of circular orbits, parameterized by energy, which then evolves adiababtically. Since energy is globally conserved, it evolves as
@@ III.4 Analytical and numerical solution on Teukolsky equation
$R^{H}_{\omega\ell m}$ $=$ $\Delta\left((r_{+}-r_{-})\frac{r-r_{+}}{r-r_{-}}\right)^{i\frac{3a-2Mr_{+}\omega}{r_{+}-r_{-}}}\,_{2}F_{1}\left(1-\ell,\ell+2;2-2iM\omega+\frac{2i\left(3a-2M^{2}\omega\right)}{r_{+}-r_{-}};\frac{r_{+}-r}{r_{+}-r_{-}}\right)\,.$ (35)
where we have introduced $\delta=\sqrt{M^{2}-a^{2}}$. To obtain the energy fluxes, one needs only replace these expressions in Eqs. (28)-(30). As stated before, the angular eigenfunctions and eigenvalues are treated using the Black Hole Perturbation Toolkit [75]. Since we are in the low frequency limit, these were expanded to first order in $\omega$.
@@ IV Results
Figure 3: EM energy flux at infinity, normalized to the Larmor result (36), for a charged particle in a circular prograde plus orbit of radius $r_{0}$ around a Kerr BH with $a=0.7M$ and $M\omega_{c}=0.1$. Solid lines are numerical results for different $(\ell,m)$ modes, dashed lines are analytical prediction (35) (we show the sum of flux in the two modes with symmetric azimuthal number ±mplus-or-$\pm m$, which contribute equally). Lowest orbital radius $r_{0}$ corresponds to ISCO, $r_{\text{ISCO}}\approx 3.13M$. The dipole ($\ell=m=1$) mode is dominant and it recovers the prediction of the Larmor formula for wide orbits with low orbital velocity (see Fig. 2). The other modes are only non negligible near ISCO. The $\ell=2$ $m=1$ mode exhibits an anti-resonance: for $r_{0}\approx 10M$ radiation to infinity is suppressed. This phenomenon is also present in higher order modes with $\ell+m$ odd.
Our results for EM radiation are summarized in Figs. 35. Fluxes at infinity, for the first few multipoles, are shown in Fig. 3 normalized to the Larmor result
@@
The radiation is predominantly dipolar; at large $r_{0}$ we find very good agreement between the Larmor result and the flux in the dipolar mode. Higher order modes can be non-negligible, but only for high frequency orbits inside $r_{0}<8M$. We can also see that the mode with $\ell=2,\,m=1$ is significantly suppressed when compared to the $\ell=m$ mode. We find this to be a general feature of modes with odd values of $\ell+m$; in addition, the dominant mode satisfies $\ell=|m|$ [49].
@@
Figure 4: EM energy flux at the horizon normalized to the Larmor formula (36). Solid lines are numerical results, dashed lines are analytical prediction (III.4). The flux is negative in all the modes we studied. The dipole dominates and higher order modes can be neglected. The $\ell=2$ $m=1$ mode exhibits an anti-resonance for $r_{0}\approx 20M$.
@@
The energy going into the horizon behaves in a similar way, with an exciting twist. Fluxes at the horizon are shown in Fig. 4, normalized to $\dot{E}^{\text{Larmor}}$, Eq. (36). As we noted already for non-spinning BHs, the energy flux into the horizon can be large, in fact much larger then the energy flux to infinity. This enhancement is apparent in Fig. (4), which shows that dipolar fluxes into the horizon can be orders of magnitude larger than that predicted by Larmor’s result.
@@
More interesting is the fact that the energy flux is negative on the horizon, a clear sign that these are superradiant modes [35], a consequence of the low orbital frequency of the circular orbits we are studying. Condition (22) for superradiance to take place is equivalent to requiring that $\Omega_{0}<\Omega_{H}\approx 0.2M^{-1}$. Indeed, this condition is met by all the orbits we are studying (see Fig. 2).
@@
Figure 5: Ratio of EM flux on the horizon to that at infinity. Solid (dashed) are numerical (low-frequency analytical) results for different $(\ell,m)$ modes. Fluxes at the horizon dominate the energy balance, leading to floating orbits. Anti-resonances of Figs. 34 are also apparent here.
@@
The impact of superradiance in the overall emission process is usually negligible: fluxes at the horizon are suppressed by a high power of the velocity of the orbiting particle [57]. However, this setup is exceptional. Radiation from isolated, charged particles around magnetized BHs can be dominated by the horizon. This is shown very clearly in Fig. 5, where we plot the ratio between the energy flux at the horizon and at infinity for different modes.
@@
The ratio between fluxes at the horizon and infinity in Fig. 5 is negative everywhere. In fact, for sufficiently large orbital radius, the energy flux on the horizon becomes larger (in absolute value) than the flux at infinity. For the dominant dipole mode, ”sufficiently large” means $r_{0}>9M$. This is closely related to the “horizon dominance effect” we found for Schwarzschild BHs [49].
From a dynamical perspective, the evolution of the particle is governed by a balance equation

### v2

@@ Abstract
We study the electromagnetic radiation reaction on a charged particle around a weakly magnetized Kerr black hole, by numerically solving the Teukolsky equation to find the energy fluxes in electromagnetic radiation at the horizon and at spatial infinity. We also employ analytical methods in the low-frequency limit, finding excellent agreement with the numerical results. For a wide range of parameters, the energy fluxes on the horizon are negative for all orbital radii: the modes are amplified through superradiance. More interestingly, the flux on the horizon is larger (in absolute value) than the flux at infinity for orbits with orbital radius $r_{0}\gtrsim 9M$floating orbits are a generic outcome of having magnetic fields around black holes. Particles on these orbits extract energy from the black hole through their radiation field; the net effect is an increase in the particles’ kinetic energy, so that they outspiral to infinity. Although in realistic accretion disks viscous forces are expected to dominate, floating orbits remain an interesting feature of the existence of ergoregions in black hole spacetimes.
@@ I Introduction
The Kerr solution to the vacuum Einstein field equations [1] describes the unique stationary and asymptotically flat black hole (BH) solution in General Relativity [2, 3, 4, 5]. It represents an isolated rotating BH, and is characterized by two parameters only – the mass and the angular momentum. This is the simplest theoretical description of any macroscopic object known to date!
In the last decade, there has been mounting experimental evidence pointing to BHs being commonplace in the Universe. Namely, we refer to the observations of gravitational waves from binary BH coalescences in the LIGO-Virgo-Kagra observatories [6, 7, 8], and of supermassive BHs at the center of galaxies [9, 10, 11]. So far, all observations are compatible with the dark compact objects being Kerr BHs described by General Relativity. The “Kerr paradigm” [12] will be tested with increasing precision in the near future with space-based [13] and other next generation detectors [14, 15, 16]. Thus, the study of the rich physics occurring in the vicinity of BHs is an active and timely topic.
Astrophysical BHs are not in isolation: instead, they are immersed in complex and dynamic systems [9, 10, 17, 18, 19, 11]. For example, BHs may be surrounded by an accretion disk of ionized matter [17, 18, 19, 20, 21, 22, 23], which can support appreciably strong magnetic fields. Indeed, recent observation show that BHs are commonly eveloped by magnetic fields of order $\sim 10^{4}$ and $\sim 10^{8}$ Gauss, for supermassive and stellar mass BHs, respectively [24, 25, 26, 27, 19]. Here, we consider a weak magnetic field, asymptotically uniform and aligned with the BH axis of rotation. We take the magnetic field weak enough that the geometry is described by the Kerr metric [28, 29]. This zero backreaction approximation is valid only if we restrict to a region of spacetime satisfying [30]
@@
Kerr BHs are prone to energy extraction mechanisms [31, 32, 33, 34, 35, 36, 37, 38, 39], which convert the BH rotational energy into kinetic energy of particles or occupation number of fields. In this paper we report an interesting effect of superradiance [35], whereby a charged particle orbiting a weakly magnetized Kerr BH can extract rotational energy from the latter and outspiral. This finding is another example of how magnetic fields can provide means of energy extraction from BHs [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. In a previous paper [49], we studied the the effect of radiation reaction in the weakly magnetized Schwarzschild system. We found a “horizon dominance effect” for a specific orbital configuration – the plus configuration. In this work, we will look at a similar orbital configuration – the prograde plus configuration – and show that the radiation emitted by the particle is now amplified through superradiance. This amplification is strong enough for the particle to experience a net energy gain and outspiral to infinity, following a so-called floating orbit. This type of orbit cannot be found in the unmagnetized Kerr system [50, 51], unless massive fundamental fields are invoked [52].
We focus on charged particles in equatorial-plane circular orbits around weakly magnetized Kerr BHs. We use the Teukolsky formalism [53, 54] to obtain the electromagnetic (EM) energy and angular momentum fluxes on the horizon and at infinity [55, 56, 57, 58]. By energy and angular momentum conservation, knowledge of the fluxes allows one to evolve the orbital parameters and study particle trajectories, as is usually done in the absence of magnetic fields [57, 58, 51, 52, 59, 60]. Although this orbital evolution is not studied in detail here, we show in Appendix A that particles in circular orbit do remain in circular orbit when driven by radiation reaction. Note that this calculation will only capture dissipative effects, even though the full self-force also includes a (much smaller) conservative counterpart [61, 62, 63].
@@
We use geometric units $G=c=1$, and a Gaussian electromagnetic unit system with $14\pi\varepsilon_{0}=\mu_{0}/4\pi=$ 1. The metric signature is (,+,+,+)(-,\,+,\,+,\,+)( - , + , + , + ), and Greek indices run from 00 to $3$.
@@ II.1 Weakly magnetized Kerr BH
Figure 1: Schematic representation of the four configurations of equatorial circular orbits in the weakly magnetized Kerr BH system. The BH is immersed in an asymptotically uniform magnetic field; both the BH angular momentum and the asymptotic magnetic field are orthogonal to the plane of this page. The charged particle (grey dot) is assumed to orbit the BH (black disk at the center) anticlockwise; the BH rotation, parametrized by $a$, is indicated by black arrows in the center; the cyclotron frequency is $\omega_{c}=qB_{0}/m$, where $B_{0}$ is the magnetic field strength at infinity, $q$ is the charge of the particle, and $m$ is its mass. The configurations are invariant under reflections about the plane of the paper. We also include a Newtonian interpretation of the difference between the plus and minus configurations: $\bm{F}_{\text{N}}$ and $\bm{F}_{\text{L}}$ are, respectively, the gravitational and Lorentz forces; the latter is centripetal (centrifugal) in minus (plus) configuration orbits.
@@
and $|a|\leq M$. The inner and outer horizons are located at the roots of $\Delta$,
@@ II.2 Circular orbit motion
$H\left(\pi_{\mu},x^{\mu}\right)=\frac{1}{2}\left[g^{rr}\pi_{r}^{2}+g^{\theta\theta}\pi_{\theta}^{2}+m^{2}\mathcal{F}\right]\,,$ (8)
@@
This system exhibits a reflection symmetry, given by the transformation $(a,\mathcal{L},\omega_{c})\to(-a,-\mathcal{L},-\omega_{c})$, and so we can always assume $\mathcal{L}>0$. Consequently, circular orbits can be divided into four different configurations:
@@
We obtained a full characterization of the two configurations present in the case of a magnetized Schwarzschild BH [49], where there is no distinction between prograde and retrograde orbits. The Kerr case is qualitatively similar [68], the only difference being the orbital frequency of particles following geodesics, which is now given by Eq. (6).
In the study of the magnetized Schwarzschild BH [49], we found a “horizon dominance effect” in plus configuration orbits: for sufficiently large $M\omega_{c}$, the energy flux on the horizon dominates the energy flux at infinity, even for arbitrarily wide circular orbits. This effect occurs for low orbital frequency $M\Omega_{0}\ll 1$, so that the radiation emitted is also of low frequency. Thus, when the BH is rotating, these modes are candidates for superradiant amplification [35]. In particular, for the case of particles in circular orbits, this effect only occurs if the particle is corrotating (cf. Eq. (22)). Although the high frequency prograde minus configuration orbits also stimulate superradiance, the amplification is not strong enough to lead to floating orbits, and so we focus exclusively on prograde plus configuration orbits.
@@
$\Omega_{0}\sim\omega_{p}\quad(r_{0}\ll r_{c})\ \ ,\qquad\Omega_{0}\sim\frac{\omega_{p}^{2}}{\omega_{c}}\quad(r_{0}\ll r_{c})\,.$ (15)
We will only study one particular set of parameters in this work, namely
@@
These are motivated by gravitational wave observations of nearly equal mass mergers [69, 70]. The value of $\omega_{c}$ can be as high as $M\omega_{c}\sim 10^{11}$ for an electron orbiting a supermassive BH, according to recent measurements of the magnetic fields surrounding the latter [24, 25, 26, 27]. We chose a more conservative value due to limited numerical precision, and also because for large values of $M\omega_{c}$ circular orbits become unstable under perturbations orthogonal to the equatorial plane [68]. Still, it is important to note that for different values of $a$ and $\omega_{c}$ we found qualitatively similar results.
We will therefore focus on prograde plus configuration orbits for the system parameters shown in Eq. (16). The orbital velocity profile is shown in Fig. 2 for radii outside the innermost circular orbit (ISCO), which we find to be
@@ III.1 The Teukolsky equation
and similarly for the source $T$, where the functions ${}_{s}S_{\omega\ell}(\theta)$ are the spin weighted spheroidal harmonics [74, 75]. Separation of variables yields two ordinary differential equations, one in the radial coordinate $r$ and another in the $\theta$ coordinate. The angular equation is a regular Sturm-Liouville problem with eigenvalues ${}_{s}E_{\omega\ell}$, and defines spin-weighted spheroidal harmonics. The radial equation is a Schrödinger-like equation; it takes the form, omitting the subscripts in $R$ and $T$ for simplicity,
@@
Thus, the problem amounts to solving the radial equation (19) with appropriate boundary conditions, corresponding to purely ingoing waves ot the horizon and purely outgoing waves at infinity. This is done by first constructing two linearly independent solutions – $R^{H}$ and $R^{\infty}$ – to the homogeneous problem ($T=0$), such that they satisfy the intended boundary conditions at the horizon and at infinity, respectively. A straightforward calculation [54] reveals that these solutions must have the following asymptotics:
@@
The solution to the non-homogeneous problem ($T\neq 0$) is then simply
@@ III.2 EM radiation by charged particle in a circular orbit
where $J_{2}$ is defined in Eq. 3.8 of Ref. [54]. The functional form of the EM 4-current $J_{\mu}$ is the same as in the case where the background is the Schwarzschild solution, and reads [49]
@@
where $\omega=m\Omega_{0}$, $\Theta(\cdot)$ is the Heaviside step function, $g_{\mu\nu}$ are the components of the Kerr metric (2), and all quantities in $Z^{\infty,H}_{\ell m}$ are evaluated at $r=r_{0}$, $\theta=\pi/2$; recall that the functions $R^{\infty,H}$ carry subscripts $\omega\ell m$.
@@
Thus, once the homogeneous radial Eq. (19) is solved, we can determine the full solution. In particular, from $R^{\infty,H}$ we can determine the energy and angular momentum fluxes in a straightforward manner. We will solve the homogeneous equation by employing both analytical and numerical methods.
@@ III.3 Comment on adiabatic evolution
Our results can be used to study the motion of a charged particle around a weakly magnetized Kerr BH, by taking a sequence of circular orbits, parameterized by energy. Since energy is globally conserved, the particle’s energy evolves as
@@ III.4 Analytical and numerical solution on Teukolsky equation
$R^{H}_{\omega\ell m}$ $=$ $\Delta\left((r_{+}-r_{-})\frac{r-r_{+}}{r-r_{-}}\right)^{i\frac{3a-2Mr_{+}\omega}{r_{+}-r_{-}}}\,_{2}F_{1}\left(1-\ell,\ell+2;2-2iM\omega+\frac{2i\left(3a-2M^{2}\omega\right)}{r_{+}-r_{-}};\frac{r_{+}-r}{r_{+}-r_{-}}\right)\,,$ (35)
where we have introduced $\delta=\sqrt{M^{2}-a^{2}}$. To obtain the energy fluxes, one needs only replace these expressions in Eqs. (28)-(30). As stated above, the angular eigenfunctions and eigenvalues are obtained using the Black Hole Perturbation Toolkit [77]. Since we are in the low frequency limit, these were expanded to first order in $\omega$.
@@ IV Results
Figure 3: EM energy flux at infinity, normalized to the Larmor result (36), for a charged particle in a circular prograde plus orbit of radius $r_{0}$ around a Kerr BH with $a=0.7M$ and $M\omega_{c}=0.1$. Solid lines are numerical results for different $(\ell,m)$ modes, dashed lines are analytical prediction (35) (we show the sum of flux in the two modes with symmetric azimuthal number ±mplus-or-$\pm m$, which contribute equally). Lowest orbital radius $r_{0}$ corresponds to ISCO, $r_{\text{ISCO}}\approx 3.13M$. The dipole ($\ell=m=1$) mode is dominant and it recovers the prediction of the Larmor formula for wide orbits with low orbital velocity (see Fig. 2). The other modes are only non negligible near ISCO. The $\ell=2,m=1$ mode exhibits an anti-resonance: for $r_{0}\approx 10M$ radiation to infinity is suppressed. This phenomenon is also present in higher order modes with $\ell+m$ odd.
Our results for EM radiation are summarized in Figs. 35. Fluxes at infinity, for the first few multipoles, are shown in Fig. 3, normalized to the Larmor result
@@
The radiation is predominantly dipolar; at large $r_{0}$ we find very good agreement between the Larmor result and the flux in the dipolar mode. Higher order modes can be non-negligible, but only for high frequency orbits inside $r_{0}<8M$. We can also see that the mode with $\ell=2,\,m=1$ is significantly suppressed when compared with the $\ell=m$ modes. We find this to be a general feature of modes with odd values of $\ell+m$; in addition, the dominant mode satisfies $\ell=|m|$ [49].
@@
Figure 4: EM energy flux at the horizon normalized to the Larmor formula (36). Solid lines are numerical results, dashed lines are analytical prediction (III.4). The flux is negative in all the modes we studied. The dipole dominates and higher order modes can be neglected. The $\ell=2,m=1$ mode exhibits an anti-resonance for $r_{0}\approx 20M$.
@@
The energy going into the horizon behaves in a similar way, with an exciting twist. Fluxes at the horizon are shown in Fig. 4, normalized to $\dot{E}^{\text{Larmor}}$, Eq. (36). As we noted already for non-spinning BHs, the energy flux into the horizon can be large, in fact much larger than the energy flux to infinity. This enhancement is apparent in Fig. (4), which shows that dipolar fluxes into the horizon can be orders of magnitude larger than those predicted by Larmor’s result.
@@
More interesting is the fact that the energy flux is negative at the horizon, a clear sign that these are superradiant modes [35], and a consequence of the low orbital frequency of the circular orbits we are studying. Condition (22) for superradiance to take place is equivalent to requiring that $\Omega_{0}<\Omega_{H}\approx 0.2M^{-1}$. Indeed, this condition is met by all the orbits we are studying (see Fig. 2).
@@
Figure 5: Ratio of EM energy flux at the horizon and at infinity. Solid (dashed) lines are numerical (low-frequency analytical) results for different $(\ell,m)$ modes. Fluxes at the horizon dominate the energy balance, leading to floating orbits. The anti-resonances of Figs. 34 are also apparent here.
@@
The impact of superradiance on the overall emission process is usually negligible: fluxes at the horizon are suppressed by a high power of the velocity of the orbiting particle [58]. However, this setup is exceptional: radiation fluxes from isolated, charged particles around magnetized BHs can be dominated by the horizon. This is shown very clearly in Fig. 5, where we plot the ratio between the energy flux at the horizon and at infinity for different modes.
Figure 6: Adiabatic evolution of orbital radius $r_{0}$, as a function of the coordinate time $t$, for a radiating charged particle around a magnetized Kerr BH. The particle has charge $q=1\,C$ and $m\approx 3.3\times 10^{-4}\,kg$, and the BH mass is $M=10M_{\odot}$. Differently colored curves correspond to different values of the cyclotron frequency $\omega_{c}$. The initial orbit is at $r_{0}=9M$, which is slightly larger (smaller) than the radius of the floating orbit for the blue (red) curve. Thus, the particle remains close to the initial orbit for a long time, eventually spiraling out (in). Larger (smaller) values of $M\omega_{c}$ than for the blue (red) curve also outspiral (inspiral). Finally, the larger the value of $\omega_{c}$ the larger the timescale of the evolution. This point explains why the blue and green curves intersect each other.
@@
The ratio between the fluxes at the horizon and at infinity in Fig. 5 is negative everywhere. In fact, for sufficiently large orbital radius, the energy flux at the horizon becomes larger (in absolute value) than the flux at infinity. For the dominant dipole mode, ”sufficiently large” means $r_{0}>9M$. This is closely related to the “horizon dominance effect” that was found for Schwarzschild BHs in [49].
From a dynamical perspective, the evolution of the particle is governed by an energy balance equation, [^2] Note this equation rests on the assumption that the charges generating the magnetic field do not exchange energy with the system.